OpenStudy (anonymous):

Find the volume of a square pyramid with base edges of 48 cm and a slant height of 26 cm.

OpenStudy (anonymous):

A. 11,520 cm3 B. 23,040 cm3 C. 7,680 cm3 D. 768 cm3

OpenStudy (anonymous):

first define the volume of a pyramid in an equation

OpenStudy (anonymous):

\[V=\frac{ 1 }{ 3 }s ^{2}h\]

OpenStudy (anonymous):

then figure out the height, do u know how to do that?

OpenStudy (anonymous):

Volume of any body = (area of base) x (height). We have the side of the base square = 48 cm, so area of base A is \[A = 48 \times 48 cm^2 = 2304 cm^2 \] From the slant height we'll find the height using Pythagoras theorem in triangle ABC \[26^2 = 24^2 + h^2\] find h and use V = A x h